OpenStudy (anonymous):

Suppose the side lengths and height of the triangle were divided by three. What effect would this have on the perimeter? & The Area? Justify your answer @Geometry

OpenStudy (anonymous):

The Orginial Perimeter was 33m and the Area was 27.

OpenStudy (anonymous):

This is the last part of the question.

OpenStudy (amistre64):

the perimeter is just like before; just divide it by 3; the area tho is nit a linear measurement and needs some work to it

OpenStudy (anonymous):

P would equal 11?

OpenStudy (anonymous):

Our base was 12? I think

OpenStudy (amistre64):

12 sounds right, youd have to dbl chk tho :)

OpenStudy (anonymous):

Our base was 12? I think

OpenStudy (anonymous):

Here let me draw it out real quick (:

OpenStudy (amistre64):

\[Area = \frac{1}{2}*base*height\] \[Area = \frac{1}{2}*\frac{base}{3}*\frac{height}{3}=\frac{1}{18}bh\]

OpenStudy (anonymous):

Our base was 12? I think

OpenStudy (anonymous):